When the quadratic equation y = (x − 2)(3x + 1) is written in standard form, what is the value of the coefficient "b"?
−5
3
−2
5

Found 2 solutions by solver91311, Shana-D77:
Answer by solver91311(24713)



Use FOIL to multiply the two binomials.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Shana-D77(132)

The standard form of a quadratic equation is y = ax^2 + bx + c.
For this type of problem, you should multiply it out (distribute) first:


y = (x − 2)(3x + 1)
y = x^2 + 1x - 6x - 2 ("FOIL")
y = x^2 - 5x - 2 (combined like terms)


Now we can see that b = -5